Instruction C++ Program Answer the question
1. Simplify A n (A U B)
2. Obtain the type of function
The function floor (x)-Z->Z
3.Evaluate the composition of the function of y=2(x+3)
4. Apply the inverse function of y=2(x+3)

Hello I found the answer Close the question please. Thank You
Answer #1

solution:

1.{4,1,0,1}

2. A n(A B}

AUA U  AUB =A (as AUA=A and A union AUB is A)

3. given function is surjective as for every y in B, there is at least one x in A such that f(x) = y

4. f(y)=2x+6

fof = f(f(y))= 2(2x+6)+6

= 4x+12+6

5.y=2(x+3)

y=2x+6

subtract 6 from both side

y-6=2x

(y-6)/2=x

f-(y)= (y-6)/2

please give me thumb up